# Ellipsoidal Density-Equalizing Map for Genus-0 Closed Surfaces

### v1

@@ 1 Introduction
Density-equalizing maps [32] are a class of mappings with shape deformation produced based on prescribed density distribution. More specifically, given a planar domain with a density distribution prescribed in it, a density-equalizing map will deform the domain with the high-density regions enlarged and the low-density regions shrunk. Over the past few decades, density-equalizing maps have been widely used in cartogram creation and data visualization [33]. Also, different variants and improved algorithms have been developed [34, 35, 36]. Choi and Rycroft [37] developed a novel method for computing surface density-equalizing maps and demonstrated its effectiveness in surface parameterization for simply connected open surfaces. Later, the method was extended for other mapping problems for simply connected open surfaces [38, 39, 40] and volumetric domains [41] with applications to medical visualization and shape modeling. Lyu et al. [42] developed a surface parameterization method for multiply connected open surfaces by combining density-equalizing maps and quasi-conformal maps. More recently, they have developed spherical density-equalizing mapping methods for genus-0 closed surfaces [43]. Using their methods, spherical area-preserving parameterizations and spherical parameterizations with controlled area change can be achieved. However, as mentioned above, mapping genus-0 closed surfaces with extreme geometry onto a spherical domain may induce large geometric distortion. For instance, the spherical area-preserving parameterization of an elongated surface may possess low area distortion but extremely high angular distortion, which is undesirable for many practical applications.
@@ 2.1 Density-equalizing maps
The above density-equalizing mapping method has been widely applied to cartogram creation and sociological data visualization. Specifically, the input density $\rho$ is commonly defined as some prescribed quantity (known as the “population”) per unit area, where the “population” can either be the actual population of a certain region on the geographical map, the average income of a region, or any other sociological data to be visualized.
In recent years, the density-equalizing mapping method has been further introduced to the field of surface parameterization. In particular, Choi et al. [37, 38] proposed methods for computing the density-equalizing maps for simply connected open surfaces in $\mathbb{R}^{3}$. Lyu et al. [42] developed parameterization methods for multiply connected open surfaces based on density-equalizing maps. More recently, Lyu et al. [43] developed a method for computing spherical density-equalizing maps for genus-0 closed surfaces.
@@ 2.2 Quasi-conformal theory
The above approximation suggests that $f$ maps an infinitesimal circle centered at $z_{0}$ to an infinitesimal ellipse centered at $f(z_{0})$. Additionally, we can determine the angles and scaling factors of the maximal magnification and maximal shrinkage using $\mu$ (see Fig. 3). More specifically, the angle of maximal magnification is given by $\operatorname{arg}(\mu(z_{0}))/2$, with the magnifying factor $|f_{z}(z_{0})|(1+|\mu(z_{0})|)$. Also, the angle of maximal shrinkage is given by $(\operatorname{arg}(\mu(z_{0}))+\pi)/2$, with the shrinking factor $|f_{z}(z_{0})(|1-|\mu(z_{0})|)$. The maximal dilation of $f$ is $K(f)=\frac{1+\|\mu\|_{\infty}}{1-\|\mu\|_{\infty}}$.
@@
Lui et al. [47] proposed an algorithm called Linear Beltrami Solver (LBS) for efficiently reconstructing a quasi-conformal mapping based on a given Beltrami coefficient $\mu$. Specifically, for any given $\mu=\rho+i\tau$, the corresponding quasi-conformal map $f=u+iv$ can be computed by solving the following equations:
@@
$\alpha_{1}=\frac{(\rho-1)^{2}+\tau^{2}}{1-\rho^{2}-\tau^{2}}\ \ \alpha_{2}=-\frac{2\tau}{1-\rho^{2}-\tau^{2}},\ \ \alpha_{3}=\frac{1+2\rho+\rho^{2}+\tau^{2}}{1-\rho^{2}-\tau^{2}}.$ (9)
@@ 3.1 Ellipsoidal density-equalizing map (EDEM)
Consider a given genus-0 closed surface $\mathcal{M}$ in $\mathbb{R}^{3}$ discretized as a triangle mesh $\left(\mathcal{V},\mathcal{E},\mathcal{F}\right)$, where $\mathcal{V}$ is the set of vertices, $\mathcal{E}$ is the set of edges, and $\mathcal{F}$ is the set of triangle faces. Our goal is to compute a bijective ellipsoidal density-equalizing map $f:\mathcal{M}\rightarrow\mathcal{E}_{a,b,c}$, where $\mathcal{E}_{a,b,c}$ is a prescribed ellipsoid with elliptic radii $a,b,c$:
@@ 3.1.1 Initial ellipsoidal parameterization
To accomplish this, we apply the Fast Ellipsoidal Conformal Mapping (FECM) method [30], which is an efficient algorithm for computing an ellipsoidal conformal parameterization by leveraging quasi-conformal theory. Initially, the input genus-0 closed surface $\mathcal{M}$ is conformally mapped to the unit sphere $\mathbb{S}^{2}$. Then, the FECM method maps the sphere onto the complex plane using the stereographic projection, followed by a Möbius transformation step that maps two desired polar points of $\mathcal{M}$ to 00 and $\infty$, respectively. However, it is important to note that the stereographic projection may lead to an uneven distribution of points. To address this issue, the FECM method incorporates an additional rescaling step to enhance the distribution of points.
@@ 3.1.2 Density-equalizing map on ellipsoid
$\mathrm{v}(\mathbf{r}(t),t)=-\frac{\nabla\rho(\mathbf{r}(t),t)}{\rho(\mathbf{r}(t),t)}.$ (12)
@@
where $[\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{r}_{k}]$ is a triangle in $N^{\mathcal{F}}(\mathbf{r}_{i})$. It is worth noting that every element of the lumped mass matrix $A_{n}$ is the area sum of the 1-ring neighborhood of the corresponding vertex at $t=t_{n}$. The matrix $L_{n}$ is a $|\mathcal{V}|\times|\mathcal{V}|$ symmetric sparse matrix with cotangent weights [48]:
@@
where $\alpha_{ij}$ and $\beta_{ij}$ are the two angles opposite to the edge $[\mathbf{r}_{i},\mathbf{r}_{j}]$. Now, Eq. (11) can be solved using the semi-discrete backward Euler method:
@@
where $\Psi^{W}_{i}(x)$, $\Psi^{W}_{j}(x)$, $\Psi^{W}_{k}(x)$ are hat functions and $\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})$, $\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{j})$, $\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{k})$ are the vertex densities at $\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{r}_{k}$. By the property $\nabla\Psi^{W}_{i}=\frac{N\times e_{ij}}{2\text{Area}(T)}$ (see Ref. [37]), the density gradient $\nabla\rho^{n+1}_{\mathcal{F}}(T)$ on the triangle $T$ is given by:
@@
where $e_{ij}$, $e_{jk}$, $e_{ki}$ are the three directed edges $[\mathbf{r}_{i}(t_{n}),\mathbf{r}_{j}(t_{n})]$, $[\mathbf{r}_{j}(t_{n}),\mathbf{r}_{k}(t_{n})]$, $[\mathbf{r}_{k}(t_{n}),\mathbf{r}_{i}(t_{n})]$, and $\mathbf{n}_{\mathcal{F}}(T)$ is the outward unit normal vector of the triangle face $T$. Then, using Eq. (23), we can obtain $\nabla\rho^{n+1}_{\mathcal{V}}$ on every vertex by:
@@
Therefore, the velocity field is given by:
$\mathrm{v}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})=-\frac{\nabla\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})}{\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})}.$ (24)
@@
$\widetilde{\mathrm{v}}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})=\mathrm{v}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})-\left(\mathrm{v}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})\cdot\mathbf{n}_{\mathcal{V}}(\mathbf{r}_{i})\right)\mathbf{n}_{\mathcal{V}}(\mathbf{r}_{i}),$ (27)
@@
$\mathbf{r}_{i}(t_{n+1})=\mathbf{r}_{i}(t_{n})-\delta t\ \widetilde{\mathrm{v}}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i}).$ (28)
@@ 3.1.3 Ensuring the bijectivity of the ellipsoidal mapping throughout the iterative process
where $a$, $b$, and $c$ are the radii of the ellipsoid. Using the rescaling transformation $h$, the initial ellipsoidal map $\mathbf{r}(0)$ and the ellipsoidal map $\mathbf{r}(t_{n})$ after the $n$-th iterative update can be mapped onto two unit spheres $\mathcal{S}^{0}:=h(\mathbf{r}(0))$ and $\mathcal{S}^{n}:=h(\mathbf{r}(t_{n}))$, respectively. It is easy to see that $h$ is a bijection and hence will not have any effect on the presence or absence of local mesh fold-overs in the ellipsoids.
@@ 3.1.5 Summary
2Compute an initial ellipsoidal conformal parameterization $f_{0}:\mathcal{M}\to\mathcal{E}_{a,b,c}$ [30];
3
4Compute the initial density $\rho^{0}_{\mathcal{F}}$ on $f_{0}(\mathcal{M})$ based on the prescribed population;
5
6Set $n=0$;
7
8repeat
9
@@ 24      Update $n=n+1$;
25
26until $\frac{\text{sd}\left(\rho^{n}_{\mathcal{V}}\right)}{\text{mean}\left(\rho^{n}_{\mathcal{V}}\right)}<\epsilon$  or  $n\geq n_{\max}$;
27The resulting bijective ellipsoidal density-equalizing map is $f:\mathcal{M}\rightarrow\mathcal{E}_{a,b,c}$ ;
@@ 3.2.3 Subproblem $E_{2}$
Here, the terms $\frac{\nabla\rho}{\rho}(\tilde{a},\tilde{b},\tilde{c})$ and $\mu(\tilde{a},\tilde{b},\tilde{c})$ are computed based on the updated ellipsoid $\mathcal{E}_{\tilde{a},\tilde{b},\tilde{c}}$. By selecting the combination with the smallest energy value, we can update the shape of the ellipsoid and reduce $E_{2}$. The shape update procedure is summarized in Algorithm 2.
@@
2Compute the energies $E_{2}(a,b,c)$, $E_{2}(a,b+\delta b,c)$, $E_{2}(a,b,c+\delta c)$, $E_{2}(a,b-\delta b,c)$, $E_{2}(a,b,c-\delta c)$, $E_{2}(a,b+\delta b,c+\delta c)$, $E_{2}(a,b+\delta b,c-\delta c)$, $E_{2}(a,b-\delta b,c+\delta c)$, $E_{2}(a,b-\delta b,c-\delta c)$;
3
4Select $(\tilde{a},\tilde{b},\tilde{c})$ that gives the smallest energy;
5
6Update the ellipsoid as $\mathcal{E}_{\tilde{a},\tilde{b},\tilde{c}}$ ;
7
@@ 3.2.4 Summary
Specifically, analogous to the EDEM algorithm, here we start by computing an initial ellipsoidal conformal parameterization (Section 3.1.1) with radii $(a,b,c)$. The initial density and the initial combined energy can be obtained using the initial ellipsoidal parameterization. Then, to solve the subproblem $E_{1}$, we deform the initial ellipsoid based on the descent direction (Section 3.2.2) iteratively. The overlap correction scheme (Section 3.1.3) and the re-coupling scheme (Section 3.1.4) are also applied at each iteration. We repeat the above process for a certain number of iterations (set by a prescribed parameter $K$) so that the energy $E_{1}$ defined based on the initial elliptic radii is sufficiently reduced. We then include an additional shape update step (Section 3.2.3) after every $K$ iteration to modify the ellipsoidal radii by solving the subproblem $E_{2}$. With the elliptic radii updated, we solve the subproblem $E_{1}$ as described above again and repeat the above process. Also, note that a finer adjustment of the elliptic radii will be needed as the iterations continue. Therefore, instead of using fixed radius step sizes $\delta b,\delta c$ in the shape update algorithm, we add a scaling factor $0.9^{m}$ to the prescribed step sizes $\delta b,\delta c$ at the $m$-th shape update step. We repeat the iterations until $|\frac{E^{n+1}-E^{n}}{E^{n}}|<\epsilon$ for some stopping parameter $\epsilon$. The final EDEQ map is obtained by $f=f_{n}$, from which we also get the final optimal elliptic radii $(a,b,c)$.
The proposed EDEQ algorithm is summarized in Algorithm 3. In practice, the initial radii $(a_{0},b_{0},c_{0})$ can be chosen arbitrarily, the time step size is set to be $\delta t=0.1$, the initial radius step sizes are set to be $\delta b=0.1$ and $\delta c=0.1$, the number of iterations for each set of fixed radii is set to be $K=5$, the error threshold is set to be $\epsilon=10^{-5}$, and the maximum number of iterations is set to be $n_{\text{max}}=300$.
@@
2Set $(a,b,c)=(a_{0},b_{0},c_{0})$;
3
4Compute an initial ellipsoidal conformal parameterization $f_{0}:\mathcal{M}\to\mathcal{E}_{a,b,c}$ [30];
5
6Compute the initial density $\rho^{0}_{\mathcal{F}}$ on $f_{0}(\mathcal{M})$ based on the prescribed population;
7
8Set $n=0$;
9
10repeat
11

### v2

@@ 1 Introduction
Density-equalizing maps [32] are a class of mappings with shape deformation produced based on a prescribed density distribution. More specifically, given a planar domain with a density distribution prescribed in it, a density-equalizing map will deform the domain with the high-density regions enlarged and the low-density regions shrunk. This concept is naturally related to the problem of cartogram creation, which focuses on creating shape deformations of geographical maps to visualize specific sociological data. In classical cartogram methods [33], continuous transformations balancing area targets with conformality were considered for districting. Over the past few decades, density-equalizing maps have been widely used in cartogram creation and data visualization [34]. Also, different variants and improved algorithms have been developed [35, 36, 37]. Choi and Rycroft [38] developed a novel method for computing surface density-equalizing maps and demonstrated its effectiveness in surface parameterization for simply connected open surfaces. Later, the method was extended to other mapping problems for simply connected open surfaces [39, 40, 41] and volumetric domains [42] with applications to medical visualization and shape modeling. Lyu et al. [43] developed a surface parameterization method for multiply connected open surfaces by combining density-equalizing maps and quasi-conformal maps. More recently, they have developed spherical density-equalizing mapping methods for genus-0 closed surfaces [44]. Using their methods, spherical area-preserving parameterizations and spherical parameterizations with controlled area change can be achieved. However, as mentioned above, mapping genus-0 closed surfaces with extreme geometry onto a spherical domain may induce large geometric distortion. For instance, the spherical area-preserving parameterization of an elongated surface may possess low area distortion but extremely high angular distortion, which is undesirable for many practical applications. Moreover, ellipsoidal models play an important role in geographic reference systems (e.g. WGS 84), and ellipsoidal mappings can be naturally utilized in cartograms and simulations of the Earth or other celestial bodies.
@@ 2.1 Density-equalizing maps
The above density-equalizing mapping method has been widely applied to cartogram creation and sociological data visualization. Specifically, the input density $\rho$ is commonly defined as some prescribed quantity (known as the “population”) per unit area, where the “population” can either be the actual population of a certain region on the geographical map, the income of a region, or any other sociological data to be visualized.
In recent years, the density-equalizing mapping method has been further introduced to the field of surface parameterization. In particular, Choi et al. [38, 39] proposed methods for computing density-equalizing maps for simply connected open surfaces in $\mathbb{R}^{3}$. Lyu et al. [43] developed parameterization methods for multiply connected open surfaces based on density-equalizing maps. More recently, Lyu et al. [44] developed a method for computing spherical density-equalizing maps for genus-0 closed surfaces.
@@ 2.2 Quasi-conformal theory
The above approximation suggests that $f$ maps an infinitesimal circle centered at $z_{0}$ to an infinitesimal ellipse centered at $f(z_{0})$. Additionally, we can determine the angles and scaling factors of the maximal magnification and maximal shrinkage using $\mu$ (see Fig. 3). More specifically, the angle of maximal magnification is given by $\operatorname{arg}(\mu(z_{0}))/2$, with the magnifying factor $|f_{z}(z_{0})|(1+|\mu(z_{0})|)$. Also, the angle of maximal shrinkage is given by $(\operatorname{arg}(\mu(z_{0}))+\pi)/2$, with the shrinking factor $|f_{z}(z_{0})|(1-|\mu(z_{0})|)$. The maximal dilation of $f$ is $K(f)=\frac{1+\|\mu\|_{\infty}}{1-\|\mu\|_{\infty}}$.
@@
Lui et al. [48] proposed an algorithm called Linear Beltrami Solver (LBS) for efficiently reconstructing a quasi-conformal mapping based on a given Beltrami coefficient $\mu$. Specifically, for any given $\mu=\xi+i\tau$, the corresponding quasi-conformal map $f=u+iv$ can be computed by solving the following equations:
@@
$\alpha_{1}=\frac{(\xi-1)^{2}+\tau^{2}}{1-\xi^{2}-\tau^{2}},\ \ \alpha_{2}=-\frac{2\tau}{1-\xi^{2}-\tau^{2}},\ \ \alpha_{3}=\frac{(\xi+1)^{2}+\tau^{2}}{1-\xi^{2}-\tau^{2}}.$ (9)
@@ 3.1 Ellipsoidal density-equalizing map (EDEM)
Consider a given genus-0 closed surface $\mathcal{M}$ in $\mathbb{R}^{3}$ discretized as a triangle mesh $\left(\mathcal{V},\mathcal{E},\mathcal{F}\right)$, where $\mathcal{V}$ is the set of vertices, $\mathcal{E}$ is the set of edges which indicate the connections between vertices and are in the form of Euclidean straight line segments in $\mathbb{R}^{3}$, and $\mathcal{F}$ is the set of triangle faces. Our goal is to compute a bijective ellipsoidal density-equalizing map $f:\mathcal{M}\rightarrow\mathcal{E}_{a,b,c}$, where $\mathcal{E}_{a,b,c}$ is a prescribed ellipsoid with elliptic radii $a,b,c$:
@@ 3.1.1 Initial ellipsoidal parameterization
To accomplish this, we apply the Fast Ellipsoidal Conformal Mapping (FECM) method [30], which is an efficient algorithm for computing an ellipsoidal conformal parameterization by leveraging quasi-conformal theory. Initially, the input genus-0 closed surface $\mathcal{M}$ is conformally mapped to the unit sphere $\mathbb{S}^{2}$. Then, the FECM method maps the sphere onto the complex plane using the stereographic projection, followed by a Möbius transformation step that maps two desired polar points of $\mathcal{M}$ to 0 and $\infty$, respectively. However, it is important to note that the stereographic projection may lead to an uneven distribution of points. To address this issue, the FECM method incorporates an additional rescaling step to enhance the distribution of points.
@@ 3.1.2 Density-equalizing map on ellipsoid
$\mathbf{v}(\mathbf{r}(t),t)=-\frac{\nabla\rho(\mathbf{r}(t),t)}{\rho(\mathbf{r}(t),t)}.$ (12)
@@
where $\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{r}_{k}]$ is a triangle in $N^{\mathcal{F}}(\mathbf{r}_{i})$. It is worth noting that every element of the lumped mass matrix $A_{n}$ is the area sum of the 1-ring neighborhood of the corresponding vertex at $t=t_{n}$. The matrix $L_{n}$ is a $|\mathcal{V}|\times|\mathcal{V}|$ symmetric sparse matrix with cotangent weights [49]:
@@
where $\alpha_{ij}$ and $\beta_{ij}$ are the two angles opposite to the edge $\mathbf{r}_{i},\mathbf{r}_{j}]$. Now, Eq. (11) can be solved using the semi-discrete backward Euler method:
@@
where $\Psi^{W}_{i}(x)$, $\Psi^{W}_{j}(x)$, $\Psi^{W}_{k}(x)$ are hat functions and $\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})$, $\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{j})$, $\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{k})$ are the vertex densities at $\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{r}_{k}$. By the property $\nabla\Psi^{W}_{i}=\frac{N\times e_{ij}}{2\text{Area}(T)}$ (see Ref. [38]), where $N$ is the unit normal vector of $T$ and $e_{ij}$ is the directed edge $\mathbf{r}_{i},\mathbf{r}_{j}]$, the density gradient $\nabla\rho^{n+1}_{\mathcal{F}}(T)$ on the triangle $T$ is given by:
@@
where $e_{ij}$, $e_{jk}$, $e_{ki}$ are the three directed edges $\mathbf{r}_{i}(t_{n}),\mathbf{r}_{j}(t_{n})]$, $\mathbf{r}_{j}(t_{n}),\mathbf{r}_{k}(t_{n})]$, $\mathbf{r}_{k}(t_{n}),\mathbf{r}_{i}(t_{n})]$, and $\mathbf{n}_{\mathcal{F}}(T)$ is the outward unit normal vector of the triangle face $T$. Then, using Eq. (14), we can obtain $\nabla\rho^{n+1}_{\mathcal{V}}$ on every vertex by:
@@
After obtaining the vertex density $\rho^{n}_{\mathcal{V}}$ and its gradient $\nabla\rho^{n}_{\mathcal{V}}$, the velocity at each vertex can be computed using Eq. (12) as follows:
$\mathbf{v}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})=-\frac{\nabla\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})}{\rho^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})}.$ (24)
@@
$\widetilde{\mathbf{v}}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})=\mathbf{v}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})-\left(\mathbf{v}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i})\cdot\mathbf{n}_{\mathcal{V}}(\mathbf{r}_{i})\right)\mathbf{n}_{\mathcal{V}}(\mathbf{r}_{i}),$ (27)
@@
$\mathbf{r}_{i}(t_{n+1})=\mathbf{r}_{i}(t_{n})-\delta t\ \widetilde{\mathbf{v}}^{n+1}_{\mathcal{V}}(\mathbf{r}_{i}).$ (28)
@@ 3.1.3 Ensuring the bijectivity of the ellipsoidal mapping throughout the iterative process
where $a$, $b$, and $c$ are the radii of the ellipsoid. Using the rescaling transformation $h$, the initial ellipsoidal map $\mathbf{r}(0)$ and the ellipsoidal map $\mathbf{r}(t_{n})$ after the $n$-th iterative update can be mapped onto two unit spheres $\mathcal{S}^{0}:=h(\mathbf{r}(0))$ and $\mathcal{S}^{n}:=h(\mathbf{r}(t_{n}))$, respectively. Note that in our work, we consider surfaces discretized in the form of triangle meshes, where the connections between vertices are represented by Euclidean straight line segments in $\mathbb{R}^{3}$. Therefore, after applying the transformation $h$ on all vertices, we can examine the bijectivity of the mapping with respect to the mesh formed by the vertices and the Euclidean straight line segments. Since $h$ is a piecewise linear anisotropic rescaling map defined on the entire triangle mesh, it is easy to see that $h$ is a bijection and hence will not have any effect on the presence or absence of local mesh fold-overs in the ellipsoids.
@@ 3.1.5 Summary
21ex
3Compute an initial ellipsoidal conformal parameterization $f_{0}:\mathcal{M}\to\mathcal{E}_{a,b,c}$ [30];
4
5Compute the initial density $\rho^{0}_{\mathcal{F}}$ on $f_{0}(\mathcal{M})$ based on the prescribed population;
6
7Set $n=0$;
8
9repeat
10
@@ 25  Update $n=n+1$;
26
27until $\frac{\text{sd}\left(\rho^{n}_{\mathcal{V}}\right)}{\text{mean}\left(\rho^{n}_{\mathcal{V}}\right)}<\epsilon$ or $n\geq n_{\max}$;
28The resulting bijective ellipsoidal density-equalizing map is $f:\mathcal{M}\rightarrow\mathcal{E}_{a,b,c}$ ;
@@ 3.2.3 Subproblem $E_{2}$
[…]
@@
21ex
3Compute the energies $E_{2}(a,b,c)$, $E_{2}(a,b+\delta b,c)$, $E_{2}(a,b,c+\delta c)$, $E_{2}(a,b-\delta b,c)$, $E_{2}(a,b,c-\delta c)$, $E_{2}(a,b+\delta b,c+\delta c)$, $E_{2}(a,b+\delta b,c-\delta c)$, $E_{2}(a,b-\delta b,c+\delta c)$, $E_{2}(a,b-\delta b,c-\delta c)$;
4
[…]
6
7Update the ellipsoid as $\mathcal{E}_{\tilde{a},\tilde{b},\tilde{c}}$ ;
8
@@ 3.2.4 Summary
Specifically, analogous to the EDEM algorithm, here we start by computing an initial ellipsoidal conformal parameterization (Section 3.1.1) with radii $(a,b,c)$. The initial density and the initial combined energy can be obtained using the initial ellipsoidal parameterization. Then, to solve the subproblem $E_{1}$, we adjust the vertex position on the current ellipsoid based on the descent direction (Section 3.2.2) iteratively. The overlap correction scheme (Section 3.1.3) and the re-coupling scheme (Section 3.1.4) are also applied at each iteration. We repeat the above process for a certain number of iterations (set by a prescribed parameter $K$) so that the energy $E_{1}$ defined based on the initial elliptic radii is sufficiently reduced. We then include an additional shape update step (Section 3.2.3) after every $K$ iteration to modify the ellipsoidal radii by solving the subproblem $E_{2}$. With the elliptic radii updated, we solve the subproblem $E_{1}$ to adjust the vertex position on the updated ellipsoid as described above again and repeat the above process. Also, note that a finer adjustment of the elliptic radii will be needed as the iterations continue. Therefore, instead of using fixed radius step sizes $\delta b,\delta c$ in the shape update algorithm, we add a scaling factor $0.9^{m}$ to the prescribed step sizes $\delta b,\delta c$ at the $m$-th shape update step. We repeat the iterations until $|\frac{E^{n+1}-E^{n}}{E^{n}}|<\epsilon$ for some stopping parameter $\epsilon$. The final EDEQ map is obtained by $f=f_{n}$, from which we also get the final optimal elliptic radii $(a,b,c)$.
The proposed EDEQ algorithm is summarized in Algorithm 3. In practice, the initial radii $(a_{0},b_{0},c_{0})$ can be chosen arbitrarily, the time step size is set to be $\delta t=0.1$, the initial radius step sizes are set to be $\delta b=0.1$ and $\delta c=0.1$, the number of iterations for each set of fixed radii is set to be $K=5$, the error threshold is set to be $\epsilon=10^{-5}$, and the maximum number of iterations is set to be $n_{\text{max}}=300$, which is sufficient for ensuring satisfactory convergence in our experiments.
@@
21ex
3Set $(a,b,c)=(a_{0},b_{0},c_{0})$;
4
5Compute an initial ellipsoidal conformal parameterization $f_{0}:\mathcal{M}\to\mathcal{E}_{a,b,c}$ [30];
6
7Compute the initial density $\rho^{0}_{\mathcal{F}}$ on $f_{0}(\mathcal{M})$ based on the prescribed population;
8
9Set $n=0$;
10
11repeat
12